OpenStudy (anonymous):

Hi! Here's my problem, I'm trying to find factors of polynomials given one of their factors. Here's the example: x^4 + 2x^2 - 8x - 16, x + 2

OpenStudy (anonymous):

@ertdre Help please?

OpenStudy (anonymous):

to find the other factor of this polynomial, do (x^4+2x^2-8x-16)/x+2

OpenStudy (anonymous):

you are trying to find what times x+2 gives you x^4+2x^2-8x-16

OpenStudy (anonymous):

Tell me what you get, ask me if you have questions

OpenStudy (anonymous):

Okay. I'll be back in a second.

OpenStudy (anonymous):

Yup, still confused. I synthetically divided with -2 and got x^3 -8 Where do I go from here?

OpenStudy (anonymous):

did you ever learn how to do polynomial division?

OpenStudy (anonymous):

Algebra Two w/ Trig is the course I'm taking right now...

OpenStudy (anonymous):

im a freshman taking bc calc, but did you ever learn how to divide a polynomia, by another equation?

OpenStudy (anonymous):

example: x^2-4 divided by x+2

OpenStudy (anonymous):

Oh, I know how to divide synthetically and how to divide long. What I don't know how to do is how to find factors through dividing, even with given a factor.

OpenStudy (anonymous):

so basically, a factor is only one half. A factor needs to be multiplied by another factor to reach the end product. Since we are given the end product, we must divide it by the factor that we are given to find the other factor.

OpenStudy (anonymous):

sorry if that made no sense

OpenStudy (anonymous):

perhaps an example will help

OpenStudy (anonymous):

say we are given the number 30, and that 5 is one of its factors

OpenStudy (anonymous):

we need to find another factor

OpenStudy (anonymous):

since 5*some number= 30, we can say 5x=30

OpenStudy (anonymous):

then divide both sides by 5, which gives us x=6 as another factor

OpenStudy (anonymous):

it is the same with this problem

OpenStudy (anonymous):

x+2 * some number = x^4+2x^2-8x-16

OpenStudy (anonymous):

so we divide both sides by x+2 to find the "some number"

OpenStudy (anonymous):

some number= (x^4+2x^2-8x-16)/(x+2)

OpenStudy (anonymous):

then since you know polynomial division, do x^4+2x^2-8x-16 divided by x+2 to find your answer

OpenStudy (anonymous):

But I already did. That's how I got x^3 - 8, and I want to know how to find more factors.

OpenStudy (anonymous):

how did you get x^3-8? That doesn't seem right